OpenStudy (anonymous):

if ln(a)=2 ln(b)=3 ln(c)=5 evaluate the following: 1) ln((a^-4)/(b^-4c^-3)) 2) (ln c^-4)(ln(a/b^3))^4

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use rule 2 listed there, then rule 3 then rule 1 for \(\bf ln\left(\cfrac{a^{-4}}{b^{-4}c^{-3}}\right)\) for \(\bf \left[\cfrac{ln(c^{-4})}{ln\left(\frac{a}{b^3}\right)}\right]^4\) distribute the exponent first then use rule 2 listed there, then rule 2 again

OpenStudy (anonymous):

Just as @jdoe0001 said follow this rule and solve the problem. If u r stuck somewhere please do come back. we are ready to help ya @homeworkgeek

OpenStudy (anonymous):

I am not too sure exactly where I'm stuck at, but I keep getting the wrong answer even after following the log rules. not really sure what im missing @heybhai

OpenStudy (anonymous):

Please elaborate the point where u are stucked @homeworkgeek

OpenStudy (anonymous):

okay so this is how I attempted the first problem: \[\ln (\frac{ a ^{-4} }{ b ^{-4} c ^{-3}})\] \[\ln (a ^{-4})-\ln (b ^{-4})+\ln(c ^{-3})\] \[-4\ln (a)+4\ln (b)-3\ln (c)\] \[-4(2)+4(3)-3(5)\] \[-8+12-15\] -11

OpenStudy (anonymous):

It will be ln(a^−4)−ln(b^−4)-ln(c^−3) Note that in third term it will be -ve sign as it has come from denominator -4ln(a)+4ln(b)+3ln(c) -4(2)+4(3)+3(5) -8+12+15 19

OpenStudy (anonymous):

Thank you! that makes a lot more sense!

OpenStudy (anonymous):

Do mind of change in sign when transposing number from numerator to denominator and vice-versa and do follow the link as @jdoe0001 suggested

OpenStudy (jdoe0001):

\(\bf ln\left(\cfrac{a^{-4}}{b^{-4}c^{-3}}\right) \\ \quad \\ \quad \\ ln(a^{-4})-[ln(b^{-4}c^{-3})]\implies ln(a^{-4})-[{\color{blue}{ ln(b^{-4})+ln(c^{-3})}}] \\ \quad \\ ln(a^{-4}){\color{blue}{ -ln(b^{-4})-ln(c^{-3})}}\implies -4ln(a)-[-4ln(b)]-[-3ln(c)] \\ \quad \\ -4ln(a)+4ln(b)+3ln(c)\)